OpenStudy (anonymous):

How many roots do the following equations have? -12x2 - 25x + 5 + x3 = 0 A. 3 B. 4 C. 5 D. 2 it's not D.

OpenStudy (anonymous):

@DatChinookGuy can you help

OpenStudy (wolf1728):

I'd say 3

OpenStudy (anonymous):

how did you get 3

OpenStudy (wolf1728):

Basically, the number of roots for an equation is equal to the highest power in the equation. So, quadratics have 2 cubic equations have 3 and so on

OpenStudy (anonymous):

okay thanks that helped alot. :)

OpenStudy (wolf1728):

u r welcome

OpenStudy (anonymous):

x⁴-5x³+11x²-25x+30 = 0      ← Check to see if +11x² can be broken up to form two                                             trinomials that share a common factor. Consider                                             x⁴-5x³ ≈ x²(x²-5x+C) ... For x²-5x+C to factor, the factors of C                                             must add to -5 ... If we try -3 & -2, that would make C = +6                                             so that 11x² would be broken up into +6x²+5x². And, notice                                             that 5 is a factor of both -25 & +30. So, C=+6 looks good. x⁴-5x³+6x² + 5x²-25x+30 = 0 x²(x²-5x+6) + 5(x²-5x+6) = 0      ← Now, factor out (x²-5x+6)               (x²+5)(x²-5x+6) = 0      ← Now, factor (x²-5x+6)              (x²+5)(x-3)(x-2) = 0 Roots are for:   i. x²+5 = 0  ⇒  x = ±i√5 ii.   x-3 = 0  ⇒  x = 3 iii.   x-2 = 0  ⇒  x = 2        ANSWER x = 2, 3, i√5, -i√5 ——————————————————————————————————————      x³-x²-13x+13 = 0   x³-x² − 13x+13 = 0 x²(x-1) - 13(x-1) = 0            ← Now, factor out (x-1)        (x²-13)(x-1) = 0 Roots are for: i. x²-13 = 0  ⇒  x = ±√(13) ii.    x-1 = 0  ⇒  x = 1        ANSWER x = 1, √(13), -√(13) *eliminate 1* ****ANSWER IS 3***** Quick example for ya in this type of math all 3 roots of our polynomial equation of degree 3 are real. Since (x−3) is a factor, then x=3 is a root. Since (4x+1) is a factor, then x=−​4​​1​​ is a root. Since (x+2) is a factor, then x=−2 is a root. x^3-x^2-13x+13 = (x-1)(x^2 - 13) = (x-1) (x-√13) (x + √13) -> roots are 1, √13, -√13 x^4-5x^3+11x^2-25x+30 = (x-2)(x^3-3x^2+5x-15) = (x-2) (x-3) (x^2 + 5) = (x-2) (x-3) (x - i√5) (x + i√5) -> roots are 2, 3, i√5, -i√5 ========================= Very simple tip I learned in early years of 8th grade math. Hope this helps :) Have a nice day. Tag me on anymore you need help with. (in conclusion answer is 3)

OpenStudy (anonymous):

thanks. can you help me with aanother question

OpenStudy (anonymous):

Find all the zeros of the equation. -4x4 - 44x2 + 3600 = 0 A. 5, -5, 6i, -6i B. 5, 6i C. 5, -5, 6i, 0 D. -5, -6i

OpenStudy (anonymous):

its not b

OpenStudy (anonymous):

Well yes we know it's definitely not B, Let f(x) = x^4 – 6x^2 – 7x – 6 Note factors of the constant term –6 are ±1, ±2, ±3 and ±6 f(−2) = 0 means x+2 is a factor of f(x) f(3) = 0 means x−3 is a factor of f(x) Do long division or synthetic division or by inspection f(x) = (x+2)(x−3)(x²+x+1) (x+2)(x−3)(x²+x+1) = 0 x = −2 or x = 3 or x²+x+1 = 0 Solving the quadratic equation, in addition to the 2 real solutions, x = ½(−1 ± i√3). Noticing the listing above I wrote, we know it's not C and D as well. What do you think the answer is? :)

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

@caitlinnicole11 correct! :)

OpenStudy (anonymous):

thanks